In a railway yard goods trains arrive that at the rate of 30 trains per day. Assuming


that the inter-arrival time follows an exponential distribution and the service time


distribution is also exponential with an average of 36 minutes, calculate the probability the yard is empty 2) the average length assuming that the line capacity of the yard is 9 trains

1
Expert's answer
2022-02-14T17:27:30-0500

Here;

"\frac{\lambda}{\mu}=\rho=0.75"

(a)Probability that the yard is empty;

"P_o=\frac{1-\rho}{1-\rho^{N+1}}"

Where N=9

"P_o=\frac{1-0.75}{1-0.75^{9+1}}=0.2649"

(b)the average length

"L=\frac{1-\rho}{1-\rho^{N+1}}\displaystyle{\sum}_{n=0}^Nn\rho^n"

"L=\frac{1-0.75}{1-0.75^{9+1}}\displaystyle{\sum}_{n=0}^9n(0.75)^n"

"L=0.28×9.58=3\ \text{trains}"